Grade 12Physical Chemistry

An ionic crystal AB crystallises in bcc lattice. If the cationic radius is 1.7 Å and anionic radius is 1.8 Å,then find out the density of crystal AB (formula mass of AB is 168.5 g/mol; )

pratheeksai28
6 Years agoGrade 12

1 Answer

Arun
6 Years ago
Density =ρ= 
A
​ 
 ×a 
3
 
Z×M 
A
​ 
 
​ 
 
Mass of FeO=56+16=72
Given,
Z=?
a=5.0 
A
˚
 
A
​ 
 =72
ρ=4 g/cm 
3
 
A
​ 
 =6.023×10 
2
 3
4= 
6.023×125×10 
−24
 
Z×72
​ 
 
Z=4
Therefore, appoximately 4 molecules of FeO are present in a unit cell.
Y=4X